# MeshAnything: Artist-Created Mesh Generation with Autoregressive Transformers

### v1

@@ 1 Introduction
In recent years, the 3D community has experienced rapid advancements, with a variety of methods developed for automatically producing high-quality 3D assets. These methods, including 3D reconstruction [44, 72, 2, 3, 34, 31], 3D generation [47, 39, 64, 41, 56, 30, 57, 69, 66], and scanning [17, 26, 27], can produce 3D assets with shape and color quality comparable to manually created ones. The success of these methods reveals the potential to replace manually created 3D models with automatically produced ones in the 3D industry, including applications in games, movies, and the metaverse, significantly reducing time and labor costs.
@@
However, this potential remains largely unrealized because the current 3D industry predominantly relies on mesh-based pipelines for their superior efficiency and controllability, while methods for producing 3D assets typically use alternative 3D representations to achieve optimal results across various scenarios. Therefore, substantial efforts [42, 16, 43, 51, 12, 53] are devoted to converting other 3D representations into meshes and have achieved some success. Meshes produced by these methods approximate the shape quality of those created by human artists, which we refer to as Artist-Created Meshes (AMs), but they still fall short in addressing the aforementioned issues.
@@
This is because all meshes produced by these methods [42, 16, 43, 51, 12, 53] exhibit significantly poorer topology quality compared to AMs. As shown in Fig. 2, these methods rely on dense faces to reconstruct 3D shapes, completely ignoring geometric characteristics. Using these meshes in the 3D industry leads to three significant problems: First, converted meshes typically contain several orders of magnitude more faces compared to AMs, leading to significant inefficiencies in storage, rendering, and simulation. Moreover, the converted meshes complicate post-processing and downstream tasks in the 3D pipeline. They significantly increase the challenge for human artists in optimizing these meshes due to their chaotic and inefficient topologies. Finally, previous methods struggle to represent sharp edges and flat surfaces, resulting in oversmoothing and bumpy artifacts as shown in Fig. 2.
In this work, we aim to solve the aforementioned issues to facilitate the application of automatically generated 3D assets in the 3D industry. As mentioned earlier, all previous methods [42, 16, 43, 51, 12, 53] extract 3D meshes with excessively dense faces in a reconstruction manner, which inherently cannot solve these issues. Therefore, we diverge from previous approaches by formulating mesh extraction as a generation problem for the first time: we teach models to generate Artist-Created Meshes (AMs) that are aligned with the given 3D assets. The meshes generated by our methods mimic the shape and topology quality of those created by human artists. Consequently, our setting, namely Shape-Conditioned AM Generation, is fundamentally free from all previous issues, enabling seamless integration of the generated results into the 3D industry pipeline.
@@
Figure 2: Comparison with Marching Cubes [42] and Remesh [4]. We apply Marching Cubes and MeshAnything to ground truth shapes and then apply remeshing to the Marching Cubes results with different voxel sizes. Existing methods extract meshes in a reconstruction manner, ignoring the geometric features of the object and producing dense meshes with poor topology. These methods fundamentally fail to capture sharp edges and flat surfaces, as shown in the zoomed-in figure.
@@
However, training such a model presents significant challenges. The first challenge is constructing the dataset, as we need paired shape conditions and Artist-Created Meshes (AMs) for model training. The shape condition must be efficiently derived from as many diverse 3D representations as possible to serve as a condition during inference. Additionally, it must have sufficient precision to accurately represent 3D shapes and be efficiently processed into features that can be injected into the model. After weighing the trade-offs, we chose point clouds due to their explicit and continuous representation, ease of derivation from most 3D representations, and the availability of mature point cloud encoders [48, 49, 75].
@@
We filter out high-quality AMs from Objaverse [19, 18] and ShapeNet [7]. When obtaining paired shape conditions, a naive approach would be to sample point clouds directly from AMs. However, this leads to poor results during inference because the sampled point clouds have excessive precision, while automatically produced 3D assets cannot provide point clouds of similar quality, causing a domain gap between training and inference. To address this issue, we intentionally corrupt the shape quality of AMs. We first extract the signed distance function from AMs [63], convert it into a coarser mesh using [42], and then sample point clouds from this coarse mesh to narrow the domain gap in shape conditions between inference and training.
@@
Following [55], we use a VQ-VAE [61] to learn a mesh vocabulary and train a decoder-only transformer [62] on this vocabulary for mesh generation. To inject shape condition, we draw inspiration from the recent success of multimodal large language models (MLLM) [68, 37], where image features encoded by pre-trained image encoders are projected into the token space of the large language models for efficient multimodal understanding. Similarly, we treat the mesh tokens obtained from the trained VQ-VAE as the language token in LLMs and use a pre-trained encoder [75] to encode the point clouds into shape features, which is later projected into the mesh token space. These shape tokens are placed at the beginning of the mesh token sequences, effectively serving as the shape conditions for next-token predictions. After predictions, these predicted mesh tokens are decoded back to meshes with the VQ-VAE decoder [55].
@@
To further enhance the quality of mesh generation, we develop a novel noise-resistant decoder for robust mesh decoding. Our observation is that as the decoder in the VQ-VAE [61] is only trained with ground truth token sequences from the encoder, it could potentially lead to a domain gap when decoding the generated token sequences. To mitigate this problem, we inject the shape condition into the VQ-VAE decoder as auxiliary information for robust decoding and fine-tune it after the VQ-VAE training. This fine-tuning process involves adding noise to the mesh token sequences to simulate possible poor-quality token sequences from the decoder-only transformer, thus making the decoder robust to such poor-quality sequences.
@@
In summary, our contributions are threefold:
@@ 2.1 Mesh Extraction
Methods for extracting meshes from 3D models are numerous and have been a subject of research for decades. Following [53], we categorize these methods into two main types: Isosurface Extraction [42, 5, 16, 6, 43, 12] and Gradient-Based Mesh Optimization [9, 23, 28, 33, 52, 36, 53].
Traditional isosurface extraction methods [42, 43, 16, 21, 32, 50, 13, 12] focus on extracting a polygonal mesh that represents the level set of a scalar function, an area that has seen extensive study in various fields. The most popular method among them is Marching Cubes [42]. It divides the space into cells, within which polygons are created to approximate the surface. Marching Cubes has been widely used for mesh extraction its robustness and simplicity. To more effectively capture sharp features, Dual Contouring [50] adopts a dual representation by extracting mesh vertices per cell and estimates vertex positions based on local isosurface details. Dual Marching Cubes [46] is another advanced approach that combines the advantages of both Marching Cubes and Dual Contouring. Recently, [13] and [12] introduce a data-driven method to determine the position of the extracted mesh based on the input field.
Transitioning to more recent developments, the advent of machine learning has ushered in new techniques for generating 3D meshes [9, 23, 28, 33, 52, 36, 53]. This line of work explores using neural networks to generate 3D meshes, where the network parameters are optimized through gradient-based methods under specific loss functions. These approaches leverage the computational power of machine learning to refine mesh quality and adaptability, making them suitable for more complex geometrical structures that traditional methods may struggle to process efficiently. [23] proposes predicting a deformable tetrahedral grid for 3D shape representation. Besides, [52] employs a differentiable Marching Tetrahedra layer for mesh extraction. Similar to [52],  [53] iteratively optimizes a 3D surface mesh by representing it as the isosurface of a scalar field.
However, these approaches fundamentally differ from ours. They ignore the characteristics of the shape and inherently cannot produce meshes with efficient topology. In contrast, Shape-Conditioned AM Generation formulates mesh extraction as a generation problem for the first time, aiming to mimic human artists in mesh extraction and thereby generating Artist-Created Meshes (AMs) with hundreds of times fewer faces.
@@
The former category is currently the mainstream research focus. Methods such as [24, 66, 69] directly generate meshes in a feed-forward manner, but because they produce dense meshes with low-quality topology similar to previous mesh extraction methods, they still encounter the same issues when applied in the 3D industry. The most popular techniques in this category are LRM-based mesh generation models [30, 66, 69], which utilize transformers to generate 3D meshes in a feed-forward manner, often conditioning on images for control.
Notably, numerous 3D generation methods [47, 59, 64, 11, 58, 71, 30, 22, 10, 38, 54, 35, 14, 15, 57, 65, 60] can also produce meshes. These methods do not directly generate meshes but instead produce 3D assets in other representations. In practice, these 3D assets often need to be converted into meshes using mesh extraction methods, establishing a connection to our approach. Similar to direct mesh generation methods, these approaches primarily focus on the generation of shape and color while ignoring the topology of the mesh, resulting in dense meshes. Consequently, they face challenges when applied to the 3D industry due to their inefficient topology.
Recently, several works have focused on the second category: generating Artist-Created Meshes(AMs) [45, 1, 55, 8]. Although our approach also focuses on AM generation, it fundamentally differs from these methods. Since they lack shape conditioning, these methods must simultaneously learn the complex 3D shape distribution—which typically alone requires extensive training [30, 57]—and the topology distribution of AMs, leading to very challenging training processes. In contrast, Shape-Conditioned AM Generation only needs to learn how to construct efficient topology for a known shape, making the learning process significantly easier. Numerous mature methods [44, 34, 30, 57, 2, 3, 66, 69] already effectively produce high-quality 3D shapes; by combining our approach with these methods, we can achieve similar outcomes to those of direct AM generation methods. This will be detailed in Sec. 3.
Among these methods, the most relevant to ours is MeshGPT [55], as we follow its architecture. [55] introduced a combination of a VQ-VAE [61] and an autoregressive transformer architecture. It first learns a mesh vocabulary with the VQ-VAE and then trains the transformer on the learned vocabulary for mesh generation. However, MeshGPT’s results are limited to several categories in ShapeNet. MeshGPT requires a training GPU hours similar to ours, but our method can generalize to unlimited categories in Objaverse. This is largely due to the difference in target complexity caused by MeshGPT needing to additionally learn the complex 3D shape distribution.
@@ 3 Shape-Conditioned AM Generation
In this section, we first introduce the formal formulation for Shape-Conditioned AM Generation and compare it with previous mesh generation settings [45, 55, 1]. We show that it can achieve the same targets as previous mesh generation methods with significantly less training effort.
Shape-Conditioned AM Generation targets to estimate a conditional distribution $p(\mathcal{M}|\mathcal{S})$. In this formula, $\mathcal{M}$ refers to the Artist-Created Mesh (AM), i.e., the mesh manually modeled by human artists. $\mathcal{S}$ refers to the 3D shape information that indicates the 3D shape to which $\mathcal{M}$ should align. The input form of $\mathcal{S}$ can be diverse, such as voxels or point clouds. Therefore, this versatility allows our method to be integrated with any 3D pipeline that outputs $\mathcal{S}$, such as 3D reconstruction [44, 34], generation [47, 30], and scanning, making these methods more efficient for the 3D industry.
Figure 3: Pipeline Overview. We introduce MeshAnything, an autoregressive transformer capable of generating Artist-Created Meshes that adhere to given 3D shapes. We sample point clouds from given 3D assets, encode them into features, and inject them into the decoder-only transformer to achieve shape-conditional mesh generation.
@@
in which $p(\mathcal{M}|\mathcal{S})$ is the focus of our shape-conditioned mesh generation. In the 3D community, numerous large models [30, 57, 69, 55] aim to estimate $p(\mathcal{S}|\mathcal{C})$ using various 3D representations and demonstrate excellent results. Besides, some single scene 3D asset production methods [44, 34, 2, 3, 47, 40, 56] can also provide samples from this distribution. By integrating our framework with these existing methods, we can leverage their capabilities to enhance our mesh generation process. This integration allows for a more resource-efficient way to estimate $p(\mathcal{M}|\mathcal{C})$, significantly reducing the complexity and resources required compared to previous methods.

### v2

@@ 1 Introduction
In recent years, the 3D community has experienced rapid advancements, with a variety of methods developed for automatically producing high-quality 3D assets. These methods, including 3D reconstruction Mildenhall et al. (2020); Yu et al. (2021); Barron et al. (2021; 2022); Kerbl et al. (2023b); Huang et al. (2024), 3D generation Poole et al. (2023); Liu et al. (2023a); Wang et al. (2023); Long et al. (2023); Sun et al. (2023); Hong et al. (2023); Tang et al. (2024); Xu et al. (2024); Wei et al. (2024), and scanning Daneshmand et al. (2018); Haleem & Javaid (2019); Haleem et al. (2022), can produce 3D assets with shape and color quality comparable to manually created ones. The success of these methods reveals the potential to replace manually created 3D models with automatically produced ones in the 3D industry, including applications in games, movies, and the metaverse, significantly reducing time and labor costs.
@@
However, this potential remains largely unrealized because the current 3D industry predominantly relies on mesh-based pipelines for their superior efficiency and controllability, while methods for producing 3D assets typically use alternative 3D representations to achieve optimal results across various scenarios. Therefore, substantial efforts Lorensen & Cline (1987); Chernyaev (1995); Lorensen & Cline (1998); Shen et al. (2021b); Chen et al. (2022); Shen et al. (2023) are devoted to converting other 3D representations into meshes and have achieved some success. Meshes produced by these methods approximate the shape quality of those created by human artists, which we refer to as Artist-Created Meshes (AMs), but they still fall short in addressing the aforementioned issues.
@@
This is because all meshes produced by these methods Lorensen & Cline (1987); Chernyaev (1995); Lorensen & Cline (1998); Shen et al. (2021b); Chen et al. (2022); Shen et al. (2023) exhibit significantly poorer topology quality compared to AMs. As shown in Fig. 2, these methods rely on dense faces to reconstruct 3D shapes, completely ignoring geometric characteristics. Using these meshes in the 3D industry leads to three significant problems: First, converted meshes typically contain several orders of magnitude more faces compared to AMs, leading to significant inefficiencies in storage, rendering, and simulation. Moreover, the converted meshes complicate post-processing and downstream tasks in the 3D pipeline. They significantly increase the challenge for human artists in optimizing these meshes due to their chaotic and inefficient topologies. Finally, previous methods struggle to represent sharp edges and flat surfaces, resulting in oversmoothing and bumpy artifacts as shown in Fig. 2.
@@
Figure 2: Comparison with Marching Cubes Lorensen & Cline (1987) and Remesh Blender Development Team (2024). We apply Marching Cubes and MeshAnything to ground truth shapes and then apply remeshing to the Marching Cubes results with different voxel sizes. Existing methods extract meshes in a reconstruction manner, ignoring the geometric features of the object and producing dense meshes with poor topology. These methods fundamentally fail to capture sharp edges and flat surfaces, as shown in the zoomed-in figure.
In this work, we aim to solve the aforementioned issues to facilitate the application of automatically generated 3D assets in the 3D industry. As mentioned earlier, all previous methods Lorensen & Cline (1987); Chernyaev (1995); Lorensen & Cline (1998); Shen et al. (2021b); Chen et al. (2022); Shen et al. (2023) extract 3D meshes with excessively dense faces in a reconstruction manner, which inherently cannot solve these issues. Therefore, we diverge from previous approaches by formulating mesh extraction as a generation problem for the first time: we teach models to generate Artist-Created Meshes (AMs) that are aligned with the given 3D assets. The meshes generated by our methods mimic the shape and topology quality of those created by human artists. Consequently, our setting, namely Shape-Conditioned AM Generation, is fundamentally free from all previous issues, enabling seamless integration of the generated results into the 3D industry pipeline.
@@
However, training such a model presents significant challenges. The first challenge is constructing the dataset, as we need paired shape conditions and Artist-Created Meshes (AMs) for model training. The shape condition must be efficiently derived from as many diverse 3D representations as possible to serve as a condition during inference. Additionally, it must have sufficient precision to accurately represent 3D shapes and be efficiently processed into features that can be injected into the model. After weighing the trade-offs, we chose point clouds due to their explicit and continuous representation, ease of derivation from most 3D representations, and the availability of mature point cloud encoders Qi et al. (2017a; b); Zhao et al. (2024).
@@
We filter out high-quality AMs from Objaverse Deitke et al. (2023b; a) and ShapeNet Chang et al. (2015). When obtaining paired shape conditions, a naive approach would be to sample point clouds directly from AMs. However, this leads to poor results during inference because the sampled point clouds have excessive precision, while automatically produced 3D assets cannot provide point clouds of similar quality, causing a domain gap between training and inference. To address this issue, we intentionally corrupt the shape quality of AMs. We first extract the signed distance function from AMs Wang et al. (2022), convert it into a coarser mesh using Lorensen & Cline (1987), and then sample point clouds from this coarse mesh to narrow the domain gap in shape conditions between inference and training.
@@
Following Siddiqui et al. (2023), we use a VQ-VAE Van Den Oord et al. (2017) to learn a mesh vocabulary and train a decoder-only transformer Vaswani et al. (2017) on this vocabulary for mesh generation. To inject shape condition, we draw inspiration from the recent success of multimodal large language models (MLLM) Wu et al. (2023); Liu et al. (2024a), where image features encoded by pre-trained image encoders are projected into the token space of the large language models for efficient multimodal understanding. Similarly, we treat the mesh tokens obtained from the trained VQ-VAE as the language token in LLMs and use a pre-trained encoder Zhao et al. (2024) to encode the point clouds into shape features, which is later projected into the mesh token space. These shape tokens are placed at the beginning of the mesh token sequences, effectively serving as the shape conditions for next-token predictions. After predictions, these predicted mesh tokens are decoded back to meshes with the VQ-VAE decoder Siddiqui et al. (2023).
@@
To further enhance the quality of mesh generation, we develop a novel noise-resistant decoder for robust mesh decoding. Our observation is that as the decoder in the VQ-VAE Van Den Oord et al. (2017) is only trained with ground truth token sequences from the encoder, it could potentially lead to a domain gap when decoding the generated token sequences. To mitigate this problem, we inject the shape condition into the VQ-VAE decoder as auxiliary information for robust decoding and fine-tune it after the VQ-VAE training. This fine-tuning process involves adding noise to the mesh token sequences to simulate possible poor-quality token sequences from the decoder-only transformer, thus making the decoder robust to such poor-quality sequences.
@@
In summary, our contributions are as follows:
@@
  • Extensive experiments demonstrate that Shape-Conditioned Mesh Generation is a more suitable setting for mesh generation, and MeshAnything significantly surpasses previous mesh generation methods.
@@ 2.1 Mesh Extraction
Methods for extracting meshes from 3D models are numerous and have been a subject of research for decades. Following Shen et al. (2023), we categorize these methods into two main types: Isosurface Extraction Lorensen & Cline (1987); Bloomenthal (1988); Chernyaev (1995); Bloomenthal & Bajaj (1997); Lorensen & Cline (1998); Chen et al. (2022) and Gradient-Based Mesh Optimization Chen et al. (2019); Gao et al. (2020); Hanocka et al. (2020); Kato et al. (2018); Shen et al. (2021a); Liao et al. (2018); Shen et al. (2023).
Traditional isosurface extraction methods Lorensen & Cline (1987; 1998); Chernyaev (1995); Doi & Koide (1991); Ju et al. (2002); Schaefer et al. (2007); Chen & Zhang (2021); Chen et al. (2022) focus on extracting a polygonal mesh that represents the level set of a scalar function, an area that has seen extensive study in various fields. The most popular method among them is Marching Cubes Lorensen & Cline (1987). It divides the space into cells, within which polygons are created to approximate the surface. Marching Cubes has been widely used for mesh extraction its robustness and simplicity. Recently, Chen & Zhang (2021) and Chen et al. (2022) introduce data-driven methods to determine the position of the extracted mesh based on the input field.
Transitioning to more recent developments, the advent of machine learning has ushered in new techniques for generating 3D meshes Chen et al. (2019); Gao et al. (2020); Hanocka et al. (2020); Kato et al. (2018); Shen et al. (2021a); Liao et al. (2018); Shen et al. (2023). This line of work explores using neural networks to generate 3D meshes, where the network parameters are optimized through gradient-based methods under specific loss functions. Shen et al. (2021a) employs a differentiable Marching Tetrahedra layer for mesh extraction. Similar to Shen et al. (2021a),  Shen et al. (2023) iteratively optimizes a 3D surface mesh by representing it as the isosurface of a scalar field.
However, these approaches fundamentally differ from ours. They ignore the characteristics of the shape and inherently cannot produce meshes with efficient topology. In contrast, MeshAnything formulates mesh extraction as a generation problem for the first time, aiming to mimic human artists in mesh extraction and thereby generating Artist-Created Meshes (AMs) with hundreds of times fewer faces.
@@
The former category is currently the mainstream research focus. Methods such as Gao et al. (2022); Wei et al. (2024); Xu et al. (2024) directly generate meshes in a feed-forward manner, but because they produce dense meshes with low-quality topology similar to previous mesh extraction methods, they still encounter the same issues when applied in the 3D industry.
Notably, numerous 3D generation methods Poole et al. (2023); Tang et al. (2023b); Wang et al. (2023); Chen et al. (2024b); Tang et al. (2023a); Yang et al. (2023); Hong et al. (2023); Fang et al. (2023); Chen et al. (2023a); Liu et al. (2024b); Shi et al. (2023); Li et al. (2023); Chen et al. (2023b; 2024c); Tang et al. (2024); Wang et al. (2024); Tochilkin et al. (2024) can also produce meshes. These methods first generate 3D assets and then convert them to dense meshes using mesh extraction methods like Lorensen & Cline (1987). Consequently, they face challenges when applied to the 3D industry due to their inefficient topology.
Recently, several works have focused on the second category: generating Artist-Created Meshes(AMs) Nash et al. (2020); Alliegro et al. (2023); Siddiqui et al. (2023); Chen et al. (2024a). Although our approach also focuses on AM generation, it fundamentally differs from these methods. Since they lack shape conditioning, these methods must simultaneously learn the complex 3D shape distribution—which typically alone requires extensive training Hong et al. (2023); Tang et al. (2024)—and the topology distribution of AMs, leading to very challenging training processes. In contrast, our methods eliminate the challenge of learning the shape distribution, allowing the model to focus on learning the topology distribution. This not only significantly reduces training costs but also enhances the model’s application value.
Among these methods, the most relevant to ours is MeshGPT Siddiqui et al. (2023), as we follow its architecture. Siddiqui et al. (2023) introduced a combination of a VQ-VAE Van Den Oord et al. (2017) and an autoregressive transformer architecture. It first learns a mesh vocabulary with the VQ-VAE and then trains the transformer on the learned vocabulary for mesh generation. However, MeshGPT’s results are limited to several categories in ShapeNet. MeshGPT requires a training GPU hours similar to ours, but our method can generalize to unlimited categories in Objaverse. As shown in Fig. 3, this is largely due to the difference in target complexity caused by MeshGPT needing to additionally learn the complex 3D shape distribution.
(a) Training Perplexity (PPL)
(b) Validation Perplexity (PPL)
Figure 3: Training and validation perplexity (PPL) for the mesh generation model under different input conditions. All models are trained with the same settings as detailed in Section 5.2. The training and validation PPL of shape-conditioned mesh generation is significantly lower than that of unconditional and image-conditioned mesh generation. This indicates that the training burden of shape-conditioned mesh generation is much lower since it avoids learning the complex 3D shape distribution.
@@ 3 Shape-Conditioned AM Generation
In this section, we first introduce the formal formulation for Shape-Conditioned AM Generation and compare it with previous mesh generation settings Nash et al. (2020); Siddiqui et al. (2023); Alliegro et al. (2023). We show that it can achieve better performance and a broader range of applications compared to the settings in previous mesh generation methods, with significantly less training effort.
Shape-Conditioned AM Generation targets to estimate a conditional distribution $p(\mathcal{M}|\mathcal{S})$. In this formula, $\mathcal{M}$ refers to the Artist-Created Mesh (AM), i.e., the mesh manually modeled by human artists. $\mathcal{S}$ refers to the 3D shape information that indicates the 3D shape to which $\mathcal{M}$ should align. The input form of $\mathcal{S}$ can be diverse, such as voxels or point clouds. Therefore, this versatility allows our method to be integrated with any 3D pipeline that outputs $\mathcal{S}$, such as 3D reconstruction Mildenhall et al. (2020); Kerbl et al. (2023b), generation Poole et al. (2023); Hong et al. (2023), and scanning, making these methods more efficient for the 3D industry.
@@
in which $p(\mathcal{M}|\mathcal{S})$ is the focus of our shape-conditioned mesh generation. As shown in Fig. 3, estimating $p(\mathcal{M}|\mathcal{S})$ is much more simpler than $p(\mathcal{M}|\mathcal{C})$, proving that our setting is much easier to train than settings in privous methods.
As for $p(\mathcal{S}|\mathcal{C})$, In the 3D community, numerous large models Team (2024); Tang et al. (2024); Xu et al. (2024); Siddiqui et al. (2023) aim to estimate using various 3D representations and demonstrate excellent results. Besides, some single scene 3D asset production methods Mildenhall et al. (2020); Kerbl et al. (2023b); Barron et al. (2021; 2022); Poole et al. (2023); Liu et al. (2023b); Sun et al. (2023) can also provide samples from this distribution. By integrating our framework with these existing methods, we can leverage their capabilities to enhance our mesh generation process. This integration allows for a more resource-efficient way to estimate $p(\mathcal{M}|\mathcal{C})$, significantly reducing the complexity and resources required compared to previous methods.
Figure 4: Pipeline Overview. We introduce MeshAnything, an autoregressive transformer capable of generating Artist-Created Meshes that adhere to given 3D shapes. During training, we inject point clouds features into a decoder-only transformer and supervise it using token sequences derived from the Artist-Created meshes. After training, MeshAnything takes point clouds sampled from various 3D representations as input and generates aligned Artist-Created meshes.